# Akide Liu1  Jing Liu1  Zizheng Pan1  Yefei He2 Gholamreza Haffari1Bohan Zhuang1, 2

### v1

@@ Abstract
A critical approach for efficiently deploying computationally demanding large language models (LLMs) is Key-Value (KV) caching. The KV cache stores key-value states of previously generated tokens, significantly reducing the need for repetitive computations and thereby lowering latency in autoregressive generation. However, the size of the KV cache grows linearly with sequence length, posing challenges for applications requiring long context input and extensive sequence generation. In this paper, we present a simple yet effective approach, called MiniCache, to compress the KV cache across layers from a novel depth perspective, significantly reducing the memory footprint for LLM inference. Our approach is based on the observation that KV cache states exhibit high similarity between the adjacent layers in the middle-to-deep portion of LLMs. To facilitate merging, we propose disentangling the states into the magnitude and direction components, interpolating the directions of the state vectors while preserving their lengths unchanged. Furthermore, we introduce a token retention strategy to keep highly distinct state pairs unmerged, thus preserving the information with minimal additional storage overhead. Our MiniCache is training-free and general, complementing existing KV cache compression strategies, such as quantization and sparsity. We conduct a comprehensive evaluation of MiniCache utilizing various models including LLaMA-2, LLaMA-3, Phi-3, Mistral, and Mixtral across multiple benchmarks, demonstrating its exceptional performance in achieving superior compression ratios and high throughput. On the ShareGPT dataset, LLaMA-2-7B with 4-bit MiniCache achieves a remarkable compression ratio of up to 5.02×5.02$\times$5.02 ×, enhances inference throughput by approximately 5×5$\times$5 ×, and reduces the memory footprint by $41\%$ compared to the FP16 full cache baseline, all while maintaining near-lossless performance.
@@ 1 Introduction
Large Language Models (LLMs), exemplified by the GPT series [1, 2, 3] and the LLaMA series [4, 5, 6], have emerged as pivotal innovations within the artificial general intelligence, notably enhancing the capabilities of natural language processing. However, these models are meticulously trained using extensive computational resources [7] and massive datasets [8], which enables them to produce text that effectively mimics human writing styles and conducts complex reasoning analysis, yet raises the challenge of efficient deployment and serving. Within the inference framework of LLMs, KV caches [9] are crucial for storing pre-computed keys and values, thus avoiding repeated calculation over the preceding context to substantially enhance the deployment efficiency of LLMs. However, the increasing demand for longer sequence lengths results in cached states becoming voluminous, leading to significant memory consumption during generation. For instance, a 175B GPT-3 model [2], with a batch size of 64 and a sequence length of 4,096 tokens (both prefilled and generated), requires approximately 1,208GB of GPU memory. This requirement is 3.45×3.45$\times$3.45 × greater than the memory used to store the model’s weights. In this context, KV cache compression is of paramount importance due to its clear benefits: 1) largely reduces the memory footprint allowing for faster generation and larger batch serving; 2) significantly lowers the cost per token, demonstrating substantial commercial benefits.
@@
Figure 1: Overview of our MiniCache strategy and example results: (a) shows the observation that the KV cache states between two adjacent layers are highly similar, particularly across the middle to deep layers. The x-axis uses index/2 to represent the similarities for each pair of layers. (b) compares the performance of MiniCache, and the mean baseline, which simply averages the KV caches of two layers, using the LLaMA-3-70B model [6] on the GSM8K dataset [10]. MiniCache, which begins merging from the half-layer depth, achieves near-lossless performance. (c) highlights the primary difference between MiniCache and previous approaches. MiniCache investigates the inter-layer redundancy of KV caches along the depth dimension of LLMs, an aspect previously overlooked by intra-layer-based methods. Here, $T$ refers to the last timestamp of pre-filling, and $T+1$ denotes to the first timestamp of decoding.
@@
Existing KV cache compression efforts can be roughly categorized into two types, namely quantization and sparsity. The quantization approaches [11, 12] propose storing the KV states in low-bit numerical values. Typically, FlexGen [13] demonstrates that 4-bit KV cache quantization can achieve lossless performance. In contrast, sparsity-driven methods aim to retain only the salient tokens while evicting the rest, either heuristically [14, 15] or adaptively [16]. Some approaches [11] explore the intersection of these two types, by assigning high-bit to salient tokens and extremely low-bit to the rest of the tokens, achieving more aggressive memory gain. Despite these innovations, existing literature merely consider the intra-layer redundancy, while neglecting another important complementary direction – the inter-layer redundancy, this motivation difference is illustrated in the Figure 1(c).
@@
  • We observe a fascinating characteristic of cross-layer KV cache states: they exhibit high similarity between adjacent layers in the middle to later stages of LLMs. Additionally, we find that not all state pairs are suitable for merging.
@@ 2 Related Work
Efficient inference for LLMs. Large Language Models (LLMs) are constrained by considerable computational and memory requirements during inference, particularly in resource-constrained environments. To mitigate these challenges, a variety of efficient inference techniques have been developed. For instance, dynamic inference methods [34, 18, 35, 36], represented by mixture-of-experts (MoE) [37, 38, 39, 40, 41], adaptively select specific sub-structures of the model during the inference process based on the input data, significantly improving inference efficiency while keeping model capacity. Techniques like Multi-Query Attention [42, 43], Kernel-driven attentions [44, 45, 46, 47], and low-rank attentions [48, 49, 50, 41] approximate the functionality of traditional attention mechanisms but with more efficient implementations. Quantization strategies [51, 52, 53, 54] involve converting the model’s weights and activations into a low bit-width format, thereby reducing memory footprint and computational intensity. Sparsification approaches [55, 56, 14, 15] eliminate unnecessary elements in both model weights and token representations, further enhancing efficiency. Some closely related works, such as MoD [17] and LayerSkips [19], considered the dynamic inference nature to ignore unimportant layers according to input; however, these methods require an additional fine-tuning process or carefully designed pre-training stages, which reduces the adaptivity of these methods. MiniCache relies on inter-layer similarity observations to perform cross-layer merging, significantly reducing memory demand.
Model merging. Merging compression involves the aggregation of a model’s parameters and activations at various granularities. This process enhances the efficiency of inference in large models and facilitates huge redundancy [57]. Linear Mode Connectivity (LMC) [58] enables the fine-tuning of models from a shared pre-trained base. Commonly, weight averaging [59] is employed as an efficient technique to perform merge compression; notably, Model Soup [60] utilizes linear averaging in this context. Advanced methods like TIES Merging [61], Model Breadcrumbs [62], and DARE [63] further enhance this process by sparsifying and combining model parameters, enabling the merging of models without sacrificing performance capabilities. Additionally, Spherical Linear intERPolation (SLERP) [64] extends beyond simple weight averaging by interpolating between model parameters. The Fisher information matrix [65] and RegMean-based methods [66] further optimize merges to produce ideal weights with respect to the $\ell_{2}$ distance to generation outputs, while maintaining the privacy of training data. However, most existing works focus on merging model parameters, with the concept of depth-wise mergeability not being thoroughly explored in prior research. MiniCache focuses on the KV cache token merging in the depth dimensional of LLMs.
@@ 3 Motivation
Figure 2: Overall of our explorations and observations : (a) shows the strong baseline by performing average merging on the KV cache. (b) shows the pairwise similarity of cache states by position between adjacent layers. (c) compares between the MiniCache, simple average, and full cach baseline on five different datasets.
@@ 3.1 Cross-Layer Redundancy in KV Cache
Prior studies have revealed the ineffectiveness of middle-to-deep layers in LLMs [20]. Thus, layer-wise early exiting in this case can effectively avoid the redundant computation with minor effect on LLM performance [67, 19]. Inspired by this, we explore a layer-wise merging of KV cache in LLMs, starting with a simple baseline by averaging similar tokens across adjacent layers. We provide our key observations as follows.
@@
Observation 2: Not all tokens are equally important to merge, a few distinct pairs require retention. Recent works [16, 15] in KV cache compression have found that keeping a few salient tokens at each layer, which contribute the majority of attention scores, could be sufficient to maintain LLM performance. In our case, we speculate that certain token pairs by position in adjacent layers also exist outlier behaviours, showing strong semantic differences that make them unsuitable for merging. Based on COQA [68] and LLaMA-2-7B [5], we investigate the similarities at the level of token pairs. As shown in Figure 2(b), we find a significant portion of token pairs share high similarities across adjacent layers. However, we observe a few outlier pairs, such as index 0 and index 15, which have large margins of difference. We consider these tokens non-mergeable due to their significant differences. We also show that merging these distinct tokens results in performance degradation, corresponding to $\gamma=0$ row in Table 2. Thus, while cross-layer merging is a promising strategy for reducing memory burdens, it must be implemented with careful consideration of token-level similarities to ensure optimal performance, as shown in Figure 2(c).
@@ 4.1 Cross-Layer Compression
Our method commences with the identification of an optimal starting layer $S$. Observations in Section 3.1 indicate that the KV cache from middle-to-deep layers exhibits explicit patterns characterized by high similarity across adjacent layers. Consequently, we select the starting layer from the middle of the LLM, specifically $S=L/2$. From this layer onward, the KV pairs are presumed to be sufficiently similar across adjacent layers to warrant their consolidation. Central to this approach is a merge function, $F$, which is designed to integrate the KV caches of consecutive layers into a single, unified cache. We define ${\bm{x}}$ as the vectorized cache state of a single token, where the superscript indicates the layer index and the subscripts $k$ and $v$ denote the keys and values, respectively. Specifically, for a pair of key/value tokens at the same position in layers $l$ and $l-1$, the merged cache is computed as
@@
Figure 3: The illustration of the proposed method MiniCache. (a) depicts the cross-layer compression process. We fetch the KV caches, from layers $l$ and ${l-1}$, and merge them into shared states via Eq. (3). Additionally, we compute the $\ell_{2}$ norm for the caches to obtain their magnitudes. Furthermore, we select unmergable tokens for retention, then store merged cache, retention tokens, and magnitudes at layer $l$ in ${\bm{C}}$. (b) demonstrates the restoration process for layers $l$ and $l-1$, which includes magnitude rescaling in Eq. (2) and retention token recovery.
@@ 4.2 KV Cache Merging and Restoration
where ${\bm{e}}$ is the directional vector. This decomposition ensures that $\frac{{\bm{e}}^{l,l-1}}{\lVert{\bm{e}}^{l,l-1}\rVert}$ is a unit vector, and allows the restored states to match the $\ell_{2}$ norm of the original states, thereby preserving the cache’s information as much as possible, the restoration is shown as Figure 3(b). For brevity, we omit the subscripts $k$ and $v$, as keys and values are decomposed in the same way. For estimating the directional component ${{\bm{e}}}^{l,l-1}$, we follow SLERP [64], which adaptively handles the interpolation, which often resembles rotation-like transformations. The choice of SLERP as the merging function is strategic, as it facilitates interpolation along the shortest path on the unit sphere between two high-dimensional vectors, thereby preserving their geometric integrity, which refers to merge operation in Figure 3(a). This is crucial for maintaining the semantic and syntactic properties of the KV caches. The formula for SLERP in our context is:
@@
where $\Omega^{l,l-1}=\arccos\left(\frac{{\bm{x}}^{l}\cdot{\bm{x}}^{l-1}}{\|{\bm{x}}^{l}\|\|{\bm{x}}^{l-1}\|}\right)$ represents the angle between vectors ${\bm{x}}^{l}$ and ${\bm{x}}^{l-1}$, and $\mathrm{sin}(\cdot)$ is the $\mathrm{sine}$ function. $t$ is an interpolation hyperparameter that adjusts the relative influence of each vector on the resulting direction, tailored to the layer depth and specific characteristics of the KV pairs. Note that when we set $t=0.5$, it will become an average merging, which we consider special cases in Eq. (A). The merged cache for each token pair is then a concatenation of the directional vector, magnitude and $\Omega^{l,l-1}$, denoting as ${\bm{c}}^{l,l-1}=[{\bm{e}}^{l,l-1},{\lVert{\bm{x}}^{l-1}\rVert},{\lVert{\bm{x}}^{l}\rVert},\Omega^{l,l-1}]$, cached components as shown in Figure 3(a). Note that apart from storing the merged directional vector, we only need to store additional token-wise magnitude and angle scalars, which is memory efficient. In this way, we achieve substantial memory efficiencies through reduced redundancy while ensuring the retention of the critical functional characteristics of the original KV pairs across transformer layers.
Unmergeable token retention. Highly distinct pairs are sensitive to merging operations, leading us to propose unmergeable token retention, as shown in Figure 3(a). Despite the high similarity between KV cache states across neighbouring layers, a few sensitive distinct pairs remain that are significantly difficult to merge and share. Aligned with previous studies, these distinct tokens carry substantial semantic meanings [15, 16]. We observe that merging sensitive tokens, which preserve layer-specific information, can lead to significant performance degradation. Therefore, it is crucial to properly disentangle the shared and unique information between adjacent layers. To address this issue, we designed a token retention strategy to selectively retain tokens that cannot be merged based on their angular distance, defined as: $d({\bm{x}}^{l},{\bm{x}}^{l-1})=\frac{1}{\pi}\Omega$. For the KV caches, the minimum and maximum angular distances are determined to identify the unmergeable tokens.
@@
Next, let ${\bm{X}}\in{\mathbb{R}}^{n\times h}$ be either the key or value cache at one attention layer, where $n$ denotes the number of tokens and $h$ is the number of hidden dimensions, and ${\bm{E}}\in{\mathbb{R}}^{n\times h}$ be the shared KV cache states. Then, for each pair of neighbouring two layers, the unmergeable tokens are selected along with the token dimension by ${\bm{R}}^{l}={\bm{X}}^{l}[{\mathbb{I}}]$, ${\bm{R}}^{l-1}={\bm{X}}^{l-1}[{\mathbb{I}}]$, then restoring to our compressed caches by $\hat{{\bm{X}}}^{l}[{\mathbb{I}}]={\bm{R}}^{l}$, $\hat{{\bm{X}}}^{l-1}[{\mathbb{I}}]={\bm{R}}^{l-1}$, as shown in Figure 3(b). Overall, we share the final cache for the two layers as ${\bm{C}}^{l,l-1}=[{\bm{E}}^{l,l-1},{\bm{R}}^{l},{\bm{R}}^{l-1},{\lVert{\bm{X}}^{l-1}\rVert},{\lVert{\bm{X}}^{l}\rVert},{\mathbb{I}}]$. This cache includes the shared KV cache states, retention of unmerged tokens, magnitude vectors for each layer, and the token-keeping index, respectively. These additional components are quite lightweight, thus compared to full-layer caches, our method remains memory-efficient, as discussed in Sec. 4.3.

### v2

@@ Abstract
A critical approach for efficiently deploying computationally demanding large language models (LLMs) is Key-Value (KV) caching. The KV cache stores key-value states of previously generated tokens, significantly reducing the need for repetitive computations and thereby lowering latency in autoregressive generation. However, the size of the KV cache grows linearly with sequence length, posing challenges for applications requiring long context input and extensive sequence generation. In this paper, we present a simple yet effective approach, called MiniCache, to compress the KV cache across layers from a novel depth perspective, significantly reducing the memory footprint for LLM inference. Our approach is based on the observation that KV cache states exhibit high similarity between the adjacent layers in the middle-to-deep portion of LLMs. To facilitate merging, we propose disentangling the states into the magnitude and direction components, interpolating the directions of the state vectors while preserving their lengths unchanged. Furthermore, we introduce a token retention strategy to keep highly distinct state pairs unmerged, thus preserving the information with minimal additional storage overhead. Our MiniCache is training-free and general, complementing existing KV cache compression strategies, such as quantization and sparsity. We conduct a comprehensive evaluation of MiniCache utilizing various models including LLaMA-2, LLaMA-3, Phi-3, Mistral, and Mixtral across multiple benchmarks, demonstrating its exceptional performance in achieving superior compression ratios and high throughput. On the ShareGPT dataset, LLaMA-2-7B with 4-bit MiniCache achieves a remarkable compression ratio of up to 5.02×5.02$\times$5.02 ×, enhances inference throughput by approximately 5×5$\times$5 ×, and reduces the memory footprint by $41\%$ compared to the FP16 full cache baseline, all while maintaining near-lossless performance. Project is available at https://minicache.vmv.re
@@ 1 Introduction
Large Language Models (LLMs), exemplified by the GPT series [1, 2, 3] and the LLaMA series [4, 5, 6], have emerged as pivotal innovations within the artificial general intelligence, significantly enhancing the capabilities of natural language processing. However, these models are meticulously trained using extensive computational resources [7] and massive datasets [8], which enables them to produce text that effectively mimics human writing styles and conducts complex reasoning analysis, yet raises the challenge of efficient deployment and serving. Within the inference framework of LLMs, KV caches [9] are crucial for storing pre-computed keys and values, thus avoiding repeated calculations over the preceding context and substantially enhancing LLMs’ deployment efficiency. However, the increasing demand for longer sequence lengths results in voluminous cached states, leading to significant memory consumption during generation. For instance, a 175B GPT-3 model [2], with a batch size of 64 and a sequence length of 4,096 tokens (both prefilled and generated), requires approximately 1,208GB of GPU memory. This requirement is 3.45×3.45$\times$3.45 × greater than the memory used to store the model’s weights. In this context, KV cache compression is of paramount importance due to its clear benefits: 1) it largely reduces the memory footprint allowing for faster generation and larger batch serving; 2) it significantly lowers the cost per token, demonstrating substantial commercial benefits.
@@
Figure 1: Overview of our MiniCache strategy and example results: (a) shows the observation that the KV cache states between two adjacent layers are highly similar, particularly across the middle to deep layers. The x-axis uses index/2 to represent the similarities for each pair of layers. (b) compares the performance of MiniCache, and the mean baseline, which simply averages the KV caches of two layers, using the LLaMA-3-70B model [6] on the GSM8K dataset [10]. MiniCache, which begins merging from the half-layer depth, achieves near-lossless performance. (c) highlights the primary difference between MiniCache and previous approaches. MiniCache investigates the inter-layer redundancy of KV caches along the depth dimension of LLMs, an aspect overlooked by intra-layer-based methods. Here, $T$ refers to the last timestamp of pre-filling, and $T+1$ des to the first timestamp of decoding.
@@
Existing KV cache compression efforts can be roughly categorized into two types, namely quantization and sparsity. The quantization approaches [11, 12] propose storing the KV states in low-bit numerical values. Typically, FlexGen [13] demonstrates that 4-bit KV cache quantization can achieve lossless performance. In contrast, sparsity-driven methods aim to retain only the salient tokens while evicting the rest, either heuristically [14, 15] or adaptively [16]. Some approaches [11] explore the intersection of these two types, by assigning high-bit to salient tokens and extremely low-bit to the rest of the tokens, achieving more aggressive memory gain. Despite these innovations, existing literature merely consider the intra-layer redundancy, while neglecting another important complementary direction – the inter-layer redundancy, as illustrated in the Figure 1(c).
@@
  • We observe a fascinating characteristic of cross-layer KV cache states: high similarity between adjacent layers in the middle to later stages of LLMs. Additionally, we find that not all state pairs are suitable for merging.
@@ 2 Related Work
Efficient inference for LLMs. Large Language Models (LLMs) are constrained by considerable computational and memory requirements during inference, particularly in resource-constrained environments. To mitigate these challenges, a variety of efficient inference techniques have been developed. For instance, dynamic inference methods [34, 18, 35, 36], represented by mixture-of-experts (MoE) [37, 38, 39, 40, 41], adaptively select specific sub-structures of the model during the inference process based on the input data, significantly improving inference efficiency while keeping model capacity. Techniques like Multi-Query Attention [42, 43], Kernel-driven attentions [44, 45, 46, 47], and low-rank attentions [48, 49, 50, 41] approximate the functionality of traditional attention mechanisms but with more efficient implementations. Quantization strategies [51, 52, 53, 54] involve converting the model’s weights and activations into a low bit-width format, thereby reducing memory footprint and computational intensity. Sparsification approaches [55, 56, 14, 15] eliminate unnecessary elements in both model weights and token representations, further enhancing efficiency. Some closely related works, such as MoD [17] and LayerSkips [19], considered the dynamic inference nature to ignore unimportant layers according to input. However, these methods require an additional fine-tuning process or carefully designed pre-training stages, which reduces the adaptivity of these methods. MiniCache relies on inter-layer similarity observations to perform cross-layer merging, significantly reducing memory demand.
Model merging. Merging compression involves the aggregation of a model’s parameters and activations at various granularities. This process enhances the efficiency of inference in large models and facilitates huge redundancy [57]. Linear Mode Connectivity (LMC) [58] enables the fine-tuning of models from a shared pre-trained base. Commonly, weight averaging [59] is employed as an efficient technique to perform merge compression. Notably, Model Soup [60] utilizes linear averaging in this context. Advanced methods like TIES Merging [61], Model Breadcrumbs [62], and DARE [63] further enhance this process by sparsifying and combining model parameters, enabling the merging of models without sacrificing performance capabilities. Additionally, Spherical Linear intERPolation (SLERP) [64] extends beyond simple weight averaging by interpolating between model parameters. The Fisher information matrix [65] and RegMean-based methods [66] further optimize merges to produce ideal weights, minimizing the $\ell_{2}$ distance to generation outputs while preserving the privacy of the training data. However, most existing works focus on merging model parameters, with the concept of depth-wise mergeability not being thoroughly explored in prior research. MiniCache focuses on the KV cache token merging in the depth dimensional of LLMs.
@@ 3 Motivation
Figure 2: Overall of our explorations and observations : (a) shows the strong baseline by performing average merging on the KV cache. (b) shows the pairwise similarity of cache states between adjacent layers. (c) compares the MiniCache, simple average, and full cache baseline across five different datasets.
@@ 3.1 Cross-Layer Redundancy in KV Cache
Prior studies have revealed the ineffectiveness of middle-to-deep layers in LLMs [20]. Thus, layer-wise early exiting in this case can effectively avoid the redundant computation with minor effect on LLM performance [67, 19]. Inspired by this, we explore a layer-wise merging of KV cache in LLMs, starting with a simple baseline by averaging full tokens across adjacent layers. We provide our key observations as follows.
@@
Observation 2: Not all tokens are equally important to merge, a few distinct pairs require retention. Recent works [16, 15] in KV cache compression have found that keeping a few salient tokens at each layer, which contribute the majority of attention scores, could be sufficient to maintain LLM performance. In our case, we speculate that certain token pairs in adjacent layers also exhibit outlier behaviours, showing strong semantic differences that make them unsuitable for merging. Based on COQA [68] and LLaMA-2-7B [5], we investigate the similarities at the level of token pairs. As shown in Figure 2(b), we find a significant portion of token pairs share high similarities across adjacent layers. However, we observe a few outlier pairs, such as indices 0 and 15, with large margins of difference. We consider these tokens non-mergeable due to their significant differences. We also show that merging these distinct tokens results in performance degradation, corresponding to $\gamma=0$ row in Table 2. Thus, while cross-layer merging is a promising strategy for reducing memory burdens, it must be implemented with careful consideration of token-level similarities to ensure optimal performance, as shown in Figure 2(c).
@@ 4.1 Cross-Layer Compression
Our method commences with the identification of an optimal starting layer $S$. Observations in Section 3.1 indicate that the KV cache from middle-to-deep layers consistently exhibits patterns of high similarity across adjacent layers. Consequently, we select the starting layer from the middle of the LLM, specifically $S=L/2$. From this layer onward, the KV pairs are assumed to be sufficiently similar across adjacent layers to warrant their consolidation. Central to this approach is a merge function, $F$, which is designed to integrate the KV caches of consecutive layers into a single, unified cache. We define ${\bm{x}}$ as the vectorized cache state of a single token, where the superscript indicates the layer index and the subscripts $k$ and $v$ denote the keys and values, respectively. Specifically, for a pair of key/value tokens at the same position in layers $l$ and $l-1$, the merged cache is computed as
@@
Figure 3: The illustration of the proposed method MiniCache. (a) depicts the cross-layer compression process. We fetch the KV caches, from layers $l$ and ${l-1}$, and merge them into shared states via Eq. (3). Additionally, we compute the $\ell_{2}$ norm for the caches to obtain their magnitudes. Furthermore, we select unmergable tokens for retention, then store merged cache, retention tokens, and magnitudes at layer $l$ in ${\bm{C}}$. (b) illustrates the restoration process for layers $l$ and $l-1$, which includes magnitude rescaling in Eq. (2) and retention token recovery.
@@ 4.2 KV Cache Merging and Restoration
where ${\bm{e}}$ is the directional vector. This decomposition ensures that $\frac{{\bm{e}}^{l,l-1}}{\lVert{\bm{e}}^{l,l-1}\rVert}$ is a unit vector, and allows the restored states to match the $\ell_{2}$ norm of the original states, thereby preserving the cache’s information as much as possible. The restoration is shown as Figure 3(b). For brevity, we omit the subscripts $k$ and $v$, as keys and values are decomposed in the same way. For estimating the directional component ${{\bm{e}}}^{l,l-1}$, we follow SLERP [64], which adaptively handles the interpolation, which often resembles rotation-like transformations. The choice of SLERP as the merging function is strategic, as it facilitates interpolation along the shortest path on the unit sphere between two high-dimensional vectors, thereby preserving their geometric integrity, which refers to merge operation in Figure 3(a). This is crucial for maintaining the semantic and syntactic properties of the KV caches. The formula for SLERP in our context is:
@@
where $\Omega^{l,l-1}=\arccos\left(\frac{{\bm{x}}^{l}\cdot{\bm{x}}^{l-1}}{\|{\bm{x}}^{l}\|\|{\bm{x}}^{l-1}\|}\right)$ represents the angle between vectors ${\bm{x}}^{l}$ and ${\bm{x}}^{l-1}$, and $\mathrm{sin}(\cdot)$ is the $\mathrm{sine}$ function. $t$ is an interpolation hyperparameter that adjusts the relative influence of each vector on the resulting direction, tailored to the layer depth and specific characteristics of the KV pairs. Note that when we set $t=0.5$, it will become an average merging along the geometry surface, which we consider special cases in Eq. (A). The merged cache for each token pair is then a concatenation of the directional vector, magnitude and $\Omega^{l,l-1}$, denoting as ${\bm{c}}^{l,l-1}=[{\bm{e}}^{l,l-1},{\lVert{\bm{x}}^{l-1}\rVert},{\lVert{\bm{x}}^{l}\rVert},\Omega^{l,l-1}]$, cached components as shown in Figure 3(a). Note that apart from storing the merged directional vector, we only need to store additional token-wise magnitude and angle scalars, which is memory efficient. In this way, we achieve substantial memory efficiencies through reduced redundancy while ensuring the retention of the critical functional characteristics of the original KV pairs across transformer layers.
Unmergeable token retention. Highly distinct pairs are sensitive to merging operations, leading us to propose unmergeable token retention, as shown in Figure 3(a). Despite the high similarity between KV cache states across neighbouring layers, a few sensitive distinct pairs remain that are significantly difficult to merge and share. Aligned with previous studies, these distinct tokens carry substantial semantic meanings [15, 16]. We observe that merging sensitive tokens, which results in the loss of layer-specific information, can lead to significant performance degradation. Therefore, it is crucial to properly disentangle the shared and unique information between adjacent layers. To address this issue, we designed a token retention strategy to selectively retain tokens that cannot be merged based on their angular distance, defined as: $d({\bm{x}}^{l},{\bm{x}}^{l-1})=\frac{1}{\pi}\Omega$. For the KV caches, the minimum and maximum angular distances are determined to identify the unmergeable tokens.
@@
Next, let ${\bm{X}}\in{\mathbb{R}}^{n\times h}$ be either the key or value cache at one attention layer, where $n$ denotes the number of tokens and $h$ is the number of hidden dimensions, and ${\bm{E}}\in{\mathbb{R}}^{n\times h}$ be the shared KV cache states. For each pair of neighbouring two layers, the unmergeable tokens are selected along with the token dimension by ${\bm{R}}^{l}={\bm{X}}^{l}[{\mathbb{I}}]$, ${\bm{R}}^{l-1}={\bm{X}}^{l-1}[{\mathbb{I}}]$, then restoring to our compressed caches by $\hat{{\bm{X}}}^{l}[{\mathbb{I}}]={\bm{R}}^{l}$, $\hat{{\bm{X}}}^{l-1}[{\mathbb{I}}]={\bm{R}}^{l-1}$, as shown in Figure 3(b). Overall, we share the final cache for the two layers as ${\bm{C}}^{l,l-1}=[{\bm{E}}^{l,l-1},{\bm{R}}^{l},{\bm{R}}^{l-1},{\lVert{\bm{X}}^{l-1}\rVert},{\lVert{\bm{X}}^{l}\rVert},{\mathbb{I}}]$. This cache includes the shared KV cache states, retention of unmerged tokens, magnitude vectors for each layer, and the token-keeping index, respectively. These additional components are quite lightweight. Thus compared to full-layer caches, our method remains memory-efficient, as discussed in Sec. 4.3.